The mole fraction of H2 S o4 in 1 molal solution is equal to

Grade:12

2 Answers

Rambhaa
6 years ago
m=(Xb x 1000) / [(Xa x Ma) m is molality and X is mole fraction where a represents solvent and b represents solute
Adarsh
6 years ago
Dear Geetha,
it is given that 1 mole of h2so4 is dissolved in 1litre of water .
and 1 litre of water contains 55.55 mole of h2o.
so mole fraction of  h2so= 1 /55.55+1= 1/56.55
Thanks and Regards
Adarsh